OpenStudy (anonymous):

Solve the system of equations algebraically. Please show the steps to your solution. 8x+2y=-2 y=-5x+1

OpenStudy (anonymous):

Substitute the expression for y from the second equation into the first equation. You can then solve for x. Then substitute that value for x into the second equation to solve for y,

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

\[8x+2\left( -5x+1 \right)=-2\] Expand, collect up like terms and solve for x. Can you do thiat?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Excellent. Now substitute x=2 into either of the given equations and solve for y. Hint: the second equation looks like it's easier to work with.

OpenStudy (phi):

osprey is saying use y=-5x+1 with x=2

OpenStudy (anonymous):

so y=-5*2+1?

OpenStudy (anonymous):

Evaluate\[y=-5\left( 2 \right)+1\]

OpenStudy (anonymous):

That's right. What do you get?

OpenStudy (anonymous):

y=-9

OpenStudy (anonymous):

Terrific. That's your answer! Well done.

OpenStudy (anonymous):

thank yo u

OpenStudy (anonymous):

The solution x=2, y=-9 works for both given equations. Pretty cool, eh? You're welcome.